### v2

# Efficient Quantification of Time-Series Prediction Error: Optimal Selection Conformal Prediction

Abstract
Uncertainty is almost ubiquitous in safety-critical autonomous systems due to dynamic environments and the integration of learning-based components. Quantifying this uncertainty–particularly for time-series predictions in multi-stage optimization–is essential for safe control and verification tasks. Conformal Prediction (CP) is a distribution-free uncertainty quantification tool with rigorous finite-sample guarantees, but its performance relies on the design of the nonconformity measure, which remains challenging for time-series data. Existing methods either overfit on small datasets, or are computationally intensive on long-time-horizon problems and/or large datasets. To overcome these issues, we propose a new parameterization of the score functions and formulate an optimization program to compute the associated parameters. The optimal parameters directly lead to norm-ball regions that constitute minimal-average-radius conformal sets. We then provide a reformulation of the underlying optimization program to enable faster computation. We provide theoretical proofs on both the validity and efficiency of predictors constructed based on the proposed approach. Numerical results on various case studies demonstrate that our method outperforms state-of-the-art methods in terms of efficiency, with much lower computational requirements.
I Introduction
Uncertainty is almost ubiquitous in safety-critical autonomous systems. The dynamic nature of external environments (e.g., autonomous driving) and the incorporation of learning-based methods (e.g., neural-networks) introduce uncertainties into the systems, which pose new challenges to safe controller design and verification. To address this issue, one way is to quantify uncertainty, in particular, for the time-series predictions that arise in multi-stage optimization problems. We classify uncertainty quantification methods into two main streams:
Bayesian methods and concentration-bounds
Uncertainty quantification using Bayesian methods include Bayesian Inference, Bayesian Neural Network and other variants [fortunato2017bayesian, neal2011mcmc, kingma2013auto]. Alternative approaches include concentrations bounds, such as Chernoff-Hoeffding (e.g., [legay2019statistical]), Clopper-Pearson (e.g., [wang2019statistical]). A more detailed review can be found in Section 2.3 of the survey paper [lindemann2024formal]. However, Bayesian methods do not have finite-sample guarantees and become computationally intractable for large-scale problems; concentration bounds can be conservative in uncertainty quantification. An alternative to these methods is conformal prediction, which are presented next.
Conformal Prediction
Conformal Prediction (CP) constitutes a statistical framework that has been introduced [vovk2005algorithmic] as a distribution-free and statistically rigorous tool to predict a $(1-\epsilon)$-confidence region for the trained prediction model, with $\epsilon\in(0,1)$. Under mild assumptions on a calibration dataset, CP provides tight finite-sample guarantees on both marginal coverage and conditional coverage. Although closely related with scenario optimization [campi2018introduction], a tool that has also been used for tight uncertainty quantification [margellos2014road], CP focuses on a different aspect and thus complements the scenario approach (see [o2025bridging] for some connections between CP and the scenario approach). Thanks to its simplicity, flexibility, and computational-efficiency, CP has been applied in probabilistic safe control synthesis and verification, such as moving-objects avoidance control [lindemann2023safe-a, stamouli2024recursively, tonkens2023scalable, yu2023signal], probabilistic reachability analysis [bortolussi2019neural, cairoli2021neural, cairoli2023conformal], probabilistic reachable sets construction [hashemi2023data, hashemi2024statistical, tebjou2023data].
However, while the confidence region produced by CP comes with rigorous probabilistic guarantees, its performance (e.g., size and shape of the region) still depends critically on one of its core components—the nonconformity measure (also called score function). Designing an appropriate score function for time-series uncertainty quantification is non-trivial. Existing works [stankeviciute2021conformal, sun2024copula, Tumu2024multi-modal, cleaveland2024conformal] either suffer from 1) overly conservative confidence regions [stankeviciute2021conformal], 2) overfitting or fitting-errors [sun2024copula, Tumu2024multi-modal], or 3) computational intractability issues [Tumu2024multi-modal, cleaveland2024conformal]. To the best of our knowledge, no work seems to alleviate these issues at the same time.
In this paper, we aim to overcome these challenges when using CP for time series, thus opening the road for its use in multi-stage optimization and safety problems that require uncertainty quantification over entire trajectories rather than single time-steps. We propose a new parameterized score function that can be optimized to provide minimal-average-radius CP regions. Our CP method generates norm-ball regions, that are convex and as we will show also tight, for multi-dimensional time series and exhibits lower computational requirements compared to other algorithmic alternatives. Our proposed approach is directly applicable to control problems such as safe learning-based MPC [lindemann2023safe-a, stamouli2024recursively, tonkens2023scalable, yu2023signal] and multi-stage safety verification [Lindemann2023Conformal-b].
Our main contributions can be summarized as:
  1. 1.
    We propose a new parameterized non-conformity measure for calibrating multi-dimensional time-series data in CP, and a mixed-integer linear programming (MILP) problem to determine optimal parameter solutions. We then provide a re-formulation of this MILP with fewer constraints to enable faster computation times.
  2. 2.
    We prove that our method is valid (concept at the core of CP); we also prove that the optimal parameters result in determining the minimum average-radius conformal set for any pre-specified normed-ball region.
  3. 3.
    We evaluate the efficacy of our approach numerically on 4 case studies and show that it produces valid conformal regions with the smallest size among baselines [stankeviciute2021conformal, sun2024copula, Tumu2024multi-modal, cleaveland2024conformal, gal2016dropout]. Specifically, the results suggest that our proposed approach reduces the conformal set size by 16.03%, 14.32%, 14.01%, 16.93% on the 4 case studies, respectively, compared to the previous State-of-the-Art (SOTA) method [cleaveland2024conformal].
  4. 4.
    Our optimization program runtime requirements are mild; compared to previous SOTA [cleaveland2024conformal], it leads to 8812.0, 78622.0, 14.4, 22.1 times faster computation on the 4 benchmark studies we have investigated numerically, respectively.
The remainder of this paper is organized as follows: Section II introduces the problem setting and the conformal prediction. In Section III we formally propose our approach, which we term Optimal Selection Conformal Prediction (OSCP). Then Section IV compares our method with 5 baseline methods via numerical experiments on 3 synthetic datasets and 1 real dataset. Finally, Section V concludes the study.
II Problem Setting and Conformal Prediction Preliminaries
II-A Problem Setting
In a discrete-time control system, let $\hat{\mathbf{Y}}_{0:T-1}=(\hat{Y}_{0},...,\hat{Y}_{T-1})\in\mathcal{Y}\subseteq\mathbb{R}^{d\times T}$ and $\mathbf{Y}_{0:T-1}=(Y_{0},...,Y_{T-1})\in\mathcal{Y}\subseteq\mathbb{R}^{d\times T}$ denote the nominal (predicted) and true trajectories of a parameter $Y_{t}$ evolving over a horizon of $T$ time-steps. For example, $\mathbf{Y}_{0:T-1}$ can be the trajectory of a moving obstacle to be avoided, while $\hat{\mathbf{Y}}_{0:T-1}$ is the predicted trajectory given by a neural-network. As another example, $\hat{\mathbf{Y}}_{0:T-1}$ can be the system state trajectory provided by the nominal system model which does not account for noise/disturbance, and thus different from the real trajectory $\mathbf{Y}_{0:T-1}$. Let $\mathbf{\tilde{Y}}_{0:T-1}:=(\tilde{Y}_{0},...,\tilde{Y}_{T-1})$ be the residual sequence capturing the error between the nominal and the true trajectory, i.e., $\tilde{Y}_{t}:=Y_{t}-\hat{Y}_{t}$.
We stipulate that the residual sequence $\mathbf{\tilde{Y}}_{0:T-1}$ is a random quantity distributed according to a probability measure $\mathbb{P}$. We assume that the corresponding probability space is defined as appropriate.
We assume throughout that we are given an exchangeable calibration dataset $D_{\mathrm{cal}}=\{\mathbf{\tilde{Y}}_{0:T-1}^{(i)}\}_{i=1}^{N}$ containing the residual sequences of historical trajectories, where the term exchangeability is defined as follows:
Definition 1 (Exchangeability).
A collection of $N$ random variables is said to be exchangeable if the joint probability distribution of any permutation of these $N$ random variables are the same. [^1] Note that exchangeability is a weaker condition compared to assuming that data are independent and identically distributed (i.i.d.).
It should be noted that we only assume that the multiple complete $T$-horizon sequences are exchangeable, without requiring exchangeability within time-horizons; i.e. we do not assume $\tilde{Y}_{t}$ and $\tilde{Y}_{t^{\prime}}$ are exchangeable or i.i.d. for a given $0\leq t\neq t^{\prime}<T$, as residuals at different time steps may exhibit temporal correlation. Another important remark is that if the nominal trajectory is generated from a data-driven model (e.g., neural-network), the calibration dataset $D_{\mathrm{cal}}$ must not involve any residual of training data, as we have assumed that all data come from the same distribution and such an operation would alter it.
For a pre-defined error level $\epsilon\in(0,1)$, our goal is to use $D_{\mathrm{cal}}$ to construct a set-value predictor $\Gamma^{\epsilon}$ that predicts a closed and bounded abstraction region $\mathcal{C}_{t}$ (such as a norm-ball) around each $\hat{Y}_{t}$ such that with at least $(1-\epsilon)$ probability, the true trajectory $\mathbf{Y}_{0:T-1}$ is completely inside these abstraction regions simultaneously for each $t=0,\ldots,T-1$.
More formally, we want to use $D_{\mathrm{cal}}$ to construct a set-valued predictor
$\Gamma^{\epsilon}:\mathbf{\hat{Y}}_{0:T-1}\mapsto\otimes_{t=0}^{T-1}\mathcal{C}_{t}\subset\mathcal{Y}$ (1)
that produces $T$ decoupled and valid (Def. 2) abstraction regions for $\mathbf{Y}_{0:T-1}$, where the term validity is defined as follows.
Definition 2 (Validity, [vovk2005algorithmic]).
Given a desired error level $\epsilon\in(0,1)$, a statistical abstraction predictor $\Gamma^{\epsilon}$ is said to be valid if for any new trajectory $\mathbf{Y}^{(\mathrm{new})}_{0:T-1}$, we have
${\mathbb{P}}\left(\mathbf{Y}^{(\mathrm{new})}_{0:T-1}\in\Gamma^{\epsilon}(\hat{\mathbf{Y}}^{(\mathrm{new})}_{0:T-1})\right)\geq 1-\epsilon,$ (2)
where $\mathbb{P}$ is the joint probability measure of the new residual sequence $\mathbf{\tilde{Y}}_{0:T-1}^{(\mathrm{new})}$ and the calibration data $D_{\mathrm{cal}}$. In the setting of this paper, (2) is equivalent to:
${\mathbb{P}}\left(Y^{(\mathrm{new})}_{t}\in\mathcal{C}_{t}\text{ for all }t=0,...,T-1\right)\geq 1-\epsilon.$ (3)
II-B Conformal Prediction and ICP Framework
Conformal Prediction (CP) is a model-agnostic and distribution-free tool that aims to quantify prediction uncertainty and produce valid prediction sets $\Gamma^{\epsilon}$ without assumptions on the data distribution. The most commonly used CP methods employ a computation-friendly framework called Inductive-Conformal-Prediction (ICP, see [vovk2005algorithmic]), sometimes also called Split-Conformal-Prediction. Such framework assumes that the data in calibration set is exchangeable with the test data. Then we use a score function (non-conformity measure) $\mathcal{A}$ to assign each calibration data with a non-conformity score $R_{i}$, and find the $p$th smallest score with $p=\lceil(1-\epsilon)(N+1)\rceil$. Then any data point with score smaller or equal to this $p$th smallest score is in the valid conformal region.
One crucial challenge is how to define the non-conformity measure $\mathcal{A}$. This is in fact non-trivial, as using a different $\mathcal{A}$ will induce CP-regions with different shapes and sizes. A CP method that has a too large CP-region is conservative and does not provide meaningful results. To evaluate the performance of a CP method, we use the term efficiency:
Definition 3 (Efficiency, [vovk2005algorithmic]).
Given an efficiency metric $\mathcal{L}_{\mathrm{eff}}$, an error level $\epsilon$, and a fixed input $\hat{\mathbf{Y}}_{0:T-1}$, a CP method $\Gamma^{\epsilon}_{1}$ is said to be more efficient than another CP method $\Gamma^{\epsilon}_{2}$ if
$\mathcal{L}_{\mathrm{eff}}\left(\Gamma^{\epsilon}_{1}(\hat{\mathbf{Y}}_{0:T-1})\right)<\mathcal{L}_{\mathrm{eff}}\left(\Gamma^{\epsilon}_{2}(\hat{\mathbf{Y}}_{0:T-1})\right).$ (4)
In the context of time-series setting where we need to produce a sequence of $T$ decoupled regions, the efficiency metric $\mathcal{L}_{\mathrm{eff}}$ is usually taken as the sum-of-widths (diameters) or sum-of-volumes (Lebesgue measure) of the $T$ CP-regions.
III Optimal Selection Conformal Prediction (OSCP)
In this paper, we propose a new CP method for time series using the Re-calibrate ICP framework [sun2024copula, Tumu2024multi-modal, cleaveland2024conformal], which splits the $D_{\mathrm{cal}}$ into 2 halves, one for learning the parameters in a parameterized score function and another for calibration (as in the standard CP). We propose a novel parameterization that provides very tight conformal regions in theory and achieves SOTA on the experiments in Section IV. We also show that the computation of this method is much more efficient than the previous SOTA method.
III-A Motivation: Minimal-Average-Radius Regions Containing $p$ Residuals
[…]
(a) 2D-regression example
(b) time-series data with $d=2,T=2$
Figure 1: Data with non-conformity scores lower or equal to $R^{[p]}$ are drawn in blue, otherwise in orange. A valid CP contains at least $p$ number of residuals inside CP regions. Motivated by simple regression case in (a), we minimize the average radius of normed-ball regions that containing p number of time-series residuals inside.
Given this fact, we now consider a simple time-series setting with $d=2$ & $T=2$ (Figure 1). At each time step, we first fix a local “center” point, and plot residual vector $\tilde{Y}_{t}^{(i)}:=Y_{t}^{(i)}-\hat{Y}_{t}^{(i)}$ with respect to this center point (think of it as the origin in x-y plane). The solid lines connecting two points denote the residuals at 2 time steps from the same time-series data, while the time series is inside the CP region if both ends of this line segment are inside the respective circles (connected via dashed lines). The first fact is that there are at least $p$ residuals of calibration data inside the CP regions (in fact there are exactly $p$ residuals when there are no “ties”). The question we seek to answer is whether we can perform a similar procedure with the single-stage regression case above. That is, construct CP-regions that give rise to norm-balls with the smallest average radius while containing at least $p$ residuals inside. We provide a positive answer to this question, and refer to our proposed method to achieve this as Optimal Selection Conformal Prediction (OSCP).
III-B OSCP: Algorithm Description, Validity and Efficiency
Considering the motivation of constructing smallest regions containing $p$ calibration data, we now present the Optimal Selection Conformal Prediction (OSCP) method using the Re-calibrate ICP framework and the Empirical-Risk-Minimization principle. This method is compatible with the common convex shapes (hyper-ball, hyper-cube, hyper-ellipsoid, etc.) for CP regions, which depends on the specific norm the user is using to calculate the absolute residual at each time step. To make the statement clearer, we first define the normed-residual-series of a residual sequence between nominal and real time-series as follows:
Definition 4 (Normed-residual-series $\hat{\boldsymbol{\epsilon}}$).
Given any norm $||\cdot||:\mathbb{R}^{d}\rightarrow\mathbb{R}$, the normed-residual-series $\tilde{\boldsymbol{e}}^{(i)}_{0:T-1}=(\tilde{e}^{(i)}_{0},...,\tilde{e}^{(i)}_{T-1})^{\top}$ for a residual sequence $\tilde{\mathbf{Y}}^{(i)}_{0:T-1}$ is defined as a time series
[…]
Using a different norm will induce different shapes of CP region for this method, which will be discussed later. Often, a natural choice is to use the $l_{2}$-norm.
Step 1: Split data
We conform to the Re-calibrate ICP framework, which requires to further split the calibration dataset. Suppose we have an exchangeable dataset $D_{\mathrm{cal}}$ drawn from $\mathbb{P}$, we split it into two disjoint subsets $D_{\mathrm{cal},1}$, $D_{\mathrm{cal},2}$ with size $n_{1}$, $n_{2}$. Although there is no requirement on how to split the dataset, in our numerical implementation we use $n_{1}\approx n_{2}$.
Step 2: Determine the optimal score function
Given a norm to calculate normed-residual-series of the calibration data, the parameterized score function $\mathcal{A}$ is defined as
$\mathcal{A}(\tilde{\mathbf{Y}}^{(i)}_{0:T-1}):=\max\left\{\tilde{e}_{0}^{(i)}-r_{0},...,\tilde{e}_{T-1}^{(i)}-r_{T-1}\right\},$ (6)
where $r_{0},...,r_{T-1}$ are parameters need to be determined. Specifically, we use $D_{\mathrm{cal},1}$ to formulate a mixed-integer linear programming problem to find optimal parameters $r_{0}^{*},...,r_{T-1}^{*}$, and the detailed formulations are in Section III-C.
Step 3: Calibrate and construct the final CP regions
Once we determine the parameters $r_{t}$, $t=0,\ldots,T-1$, based on $D_{\mathrm{cal},1}$, we have a well-defined score function $\mathcal{A}$, and as a result we can calculate non-conformity scores $R_{i}=\mathcal{A}(\tilde{\mathbf{Y}}^{(i)}_{0:T-1})$ for each data in $D_{\mathrm{cal},2}$.
Now, let $p_{2}=\lceil(1-\epsilon)(n_{2}+1)\rceil$. Suppose $R^{[p_{2}]}$ is the $p_{2}$th smallest non-conformity score among scores of $D_{\mathrm{cal},2}$, we can then construct the final CP regions $\Gamma^{\epsilon}:\hat{\mathbf{Y}}_{0:T-1}\mapsto\otimes_{t=0}^{T-1}\mathcal{C}_{t}$, where the region at each time step is:
[…]
Then these regions are valid conformal regions.
Theorem 1 (Validity).
Suppose $\mathcal{C}_{0},...,\mathcal{C}_{T-1}$ are derived via the procedures above, then we have ${\mathbb{P}}\left(Y^{(\mathrm{new})}_{t}\in\mathcal{C}_{t},\forall t=0,...,T-1\right)\geq 1-\epsilon,$ where $\mathbb{P}$ is the joint probability measure of $D_{\mathrm{cal},2}$ and $\tilde{\mathbf{Y}}_{0:T-1}^{(\mathrm{new})}$.
Proof: The construction of score function $\mathcal{A}$ depends only on $D_{\mathrm{cal},1}$, and does not include information from $D_{\mathrm{cal},2}$. Thus, the non-conformity score for each data in $D_{\mathrm{cal},2}$ is exchangeable with that of a new data drawn from $\mathcal{D}$. Let $p_{2}=\lceil(1-\epsilon)(n_{2}+1)\rceil$, by Lemma 1 in [Tibshirani2019Conformal], the conformal regions defined as $\Gamma^{\epsilon}(\hat{\mathbf{Y}}_{0:T-1}^{(\mathrm{new})}):=\left\{\mathbf{Y}\in\mathbb{R}^{d\times T}\mid\mathcal{A}(\mathbf{Y}-\hat{\mathbf{Y}}^{(\mathrm{new})}_{0:T-1})\leq R^{[p_{2}]}\right\}$ has property that ${\mathbb{P}}\left(\mathbf{Y}^{(\mathrm{new})}_{0:T-1}\in\Gamma^{\epsilon}(\hat{\mathbf{Y}}^{(\mathrm{new})}_{0:T-1})\right)\geq 1-\epsilon$. Now, for any $\mathbf{Y}\in\mathbb{R}^{d\times T}$,
[…]
Efficiency of this method
This method produces the smallest average radius (over T regions) regions that a valid CP method can achieve based on $D_{\mathrm{cal},1}$, with respect to a user’s predefined norm for calculating normed-residual-series, e.g., if a user uses $l_{2}$-norm to calculate absolute residuals, then this method produces the 2-norm balls with the minimum average radius any CP method can achieve. We formalize this in the theorem below.
Theorem 2 (Empirical Average Radius Minimization).
Suppose $\{r_{0}^{*},\dots,r_{T-1}^{*}\}$ are the optimal parameters computed in Step 2 (see also the optimization program in III-C), based on $D_{\mathrm{cal},1}$. The minimum average radius of an empirical CP-region is then equal to $\frac{1}{T}\sum_{t=0}^{T-1}r_{t}^{*}$.
Note that the term “empirical” refers to the Empirical-Risk-Minimization (ERM) principle. The regions generated by calibrating data in $D_{\mathrm{cal},1}$ via score function $\mathcal{A}(\tilde{\mathbf{Y}}^{(i)}_{0:T-1}):=\max\left\{\tilde{e}_{0}^{(i)}-r_{0}^{*},...,\tilde{e}_{T-1}^{(i)}-r_{T-1}^{*}\right\}$ are not valid CP regions, but this Theorem shows that OSCP is efficient in the sense of ERM principle. The proof can be found in Appendix VI.
Shapes of CP regions
The shape of CP regions that this method produces depends on the user pre-defined norm for calculating the absolute residuals. This can be viewed as a hyper-parameter for the method. For instance, $l_{2}$-norm produces ball-shaped regions, $l_{1}$ or $l_{\infty}$-norm induces hyper-rectangle regions, and positive-definite matrix $A$-norm results in ellipsoid-shaped regions (see [Xu2024Conformal]). Ellipsoidal regions is flexible, but may lead to over-fitting when the data is not enough to reflect its shape in high dimensional spaces. For the case that the dataset is small and we don’t have prior knowledge of the data distribution, we can assume the prediction residual follows a gaussian error and employ $l_{2}$-norm.
III-C Optimal Parameter Computation
Suppose we have select the shape of CP region by choosing a specific norm $||\cdot||$, and we have calculated the normed-residual-series for data in $D_{\mathrm{cal},1}$. Let $p_{1}=\lceil(1-\epsilon)(n_{1}+1)\rceil$. Our goal is to determine optimal parameters $r_{0},...,r_{T-1}$ by using the first calibration dataset $D_{cal,1}$.
Recalling the motivating example in III-A, we seek to determine a series of norm-balls with radii $r_{t}$, $t=0,\ldots,T-1$, that have the minimum average radius (equivalently radius sum, i.e., $\sum_{t=0}^{T-1}r_{t}$) and contain at least $p_{1}$ normed-residual-series $\tilde{\boldsymbol{e}}^{(i)}_{0:T-1}$’s. We can achieve this by means of the following optimization problem:
$\underset{\{r_{t}\},\{b_{i}\}}{\min}$ $\sum_{t=0}^{T-1}r_{t}$ (MILP)
subject to $b_{i}\cdot(\tilde{e}_{t}^{(i)}-r_{t})\leq 0,\quad t=0,\dots,T-1,$ (8)
$\qquad\qquad\qquad\qquad\ \ \ i=1,\dots,n_{1}$
$\sum_{i=1}^{n_{1}}b_{i}=p_{1}$ (9)
$b_{i}\in\{0,1\},\quad i=1,...,n_{1}.$ (10)
This is a mixed-integer linear programming problem that is always feasible, and in fact we can remove a large fraction of redundant constraints to enable faster computation while keeping the optimal solutions of $r_{0},...,r_{T-1}$ unchanged. To reduce the size of this program and improve the associated computational efficiency, two redundant constraint sets, can be identified and removed. To this end, suppose that for each t, $\tilde{e}_{t}^{[p_{1}]}$ is the $p_{1}$th smallest value among $\tilde{e}_{t}^{(1)},...,\tilde{e}_{t}^{(n_{1})}$. Then the first redundant constraint set is defined by
$S_{1}:=\left\{i\in\{1,\ldots,n_{1}\}\mid\tilde{e}_{t}^{(i)}\leq\tilde{e}_{t}^{[p_{1}]},\ \forall t=0,...,T-1\right\}.$ (11)
This set denotes the indices of all inactive constraints. In the case we are provided (or often it is easy to identify) a feasible solution $\{r_{0}^{(\mathrm{feas})},...,r_{T-1}^{(\mathrm{feas})},b_{1}^{(\mathrm{feas})},...,b_{n_{1}}^{(\mathrm{feas})}\}$ to the (MILP), then we can neglect a second redundant set
$S_{2}:=\left\{i\in\{1,...,n_{1}\}\mid\tilde{e}_{t}^{(i)}>r_{t}^{(\mathrm{feas})}\ \text{for}\ \forall t=0,...,T-1\right\},$ (12)
which includes all solutions that would lead to a cost (sum of radii) greater than that of the available feasible solution.
Although the method to find such feasible solutions is not unique, one fast and easy-to-implement heuristic procedure is as follows: for each $i=1,...,n_{1}$, we calculate the sum of normed residuals $\mathrm{TotalRes}(i):=\sum_{t=0}^{T-1}\tilde{e}^{(i)}_{t}$ and sort $\mathrm{TotalRes}(i)$’s in non-decreasing order. Then we pick the first $p_{1}$ indices, $i_{1},...,i_{p_{1}}$ and let $b_{i}^{(\mathrm{feas})}=1$ for these indices, $b_{i}^{(\mathrm{feas})}=0$ otherwise. Let $r_{t}^{(\mathrm{feas})}=\underset{i=i_{1},...i_{p_{1}}}{\max}\tilde{e}_{t}^{(i)}$. Then we have a feasible solution to (MILP).
Once $S_{1}$ & $S_{2}$ are identified, we can set up a modified optimization program as follows.
$\underset{r_{t},b_{i}}{\min}$ $\sum_{t=0}^{T-1}r_{t}$ (MILP-fast)
subject to $b_{i}\cdot(\tilde{e}_{t}^{(i)}-r_{t})\leq 0,\quad t=0,\dots,T-1,\ i\in S$ (13)
$\underset{i\in S_{1}}{\max}\{\tilde{e}_{t}^{(i)}\}\leq r_{t},\quad t=0,\dots,T-1$ (14)
$\sum_{i\in S}b_{i}=p_{1}-|S_{1}|$ (15)
$b_{i}\in\{0,1\},\quad i\in S$ (16)
$\text{where }S:=\{1,2,...,n_{1}\}\setminus(S_{1}\cup S_{2}).$
Theorem 3 (Equivalence of (MILP) & (MILP-fast)).
When $|S_{1}|<p_{1}$, (MILP-fast) is always feasible, and its set of optimal solutions of $\{r_{0}^{*},...,r_{T-1}^{*}\}$ coincides with that of (MILP). Otherwise, if $|S_{1}|\geq p_{1}$, the optimal solution to (MILP) is $r_{t}=\tilde{e}_{t}^{[p_{1}]},\ t=0,...,T-1$.
The corresponding proof is in Appendix VII. Theorem 3 implies that when $|S_{1}|<p_{1}$, we can solve (MILP-fast) to find optimal parameters $r_{0}^{*},...,r_{T-1}^{*}$. The rough idea is that to make the choice of $r_{t}$’s optimal, we must always consider containing residual-time-series from $S_{1}$ but not considering containing those from $S_{2}$. Thus, when $S_{1}<p_{1}$, solving (MILP) is equivalent to solve (MILP-fast). On the other hand, when $|S_{1}|\geq p_{1}$ (although not common in practice), there are already more than $p_{1}$ residual-time-series to be contained inside the norm-balls, so we can simply choose $r_{t}=\tilde{e}_{t}^{[p_{1}]}$, $t=0,\ldots,T-1$, as the optimal parameters.
When the error level $\epsilon$ is small, (MILP-fast) usually can remove a large number of mixed-integer constraints in (8) and integer variables $b_{i}$’s, which makes the computation much faster. The detailed results of increased running speed can be seen in Section IV-C.
The pseudo-code of this faster algorithm for computing optimal parameters is in Appendix VIII.
IV Numerical Experiments
To demonstrate the performance of our method, we test on both simulated and real time-series with different time horizons $T$, dimensions $d$, and calibration dataset sizes $N$, taken from [sun2024copula]. We compare our method with 5 baseline uncertainty-quantification (UQ) methods, and the results show that among all valid alternatives, the efficiency of our proposed approach outperforms the state-of-the-art method on all case studies.
Baseline Approaches
We selected MC-dropout [gal2016dropout] and CF-RNN [stankeviciute2021conformal] as the baseline approaches for Bayesian UQ method and CP for time series, respectively; as well as three recent approaches in CP for time series, namely, CopulaCPTS [sun2024copula] and LCP [cleaveland2024conformal] as parameter optimization methodologies, and CRD [Tumu2024multi-modal] as a convex CP baseline. Since both CRD and our method allow users to specify shapes, we test hyper-rectangle & ellipsoid shapes of CRD (denoted as CRD-Rect & CRD-Ell, respectively) and ball & ellipsoid shapes of our method ($l_{2} & OSCP$-Ell).
IV-A Synthetic Datasets
Particle trajectory
According to [sun2024copula], the first two datasets are generated from the Interacting Particle System [Kipf2018Neural], and extra Gaussian noises with standard deviations $\sigma=0.01$ and $0.05$ are added to the dynamics of particle simulations in two datasets, respectively. For each case, data is split into 2000/500/500 for training, calibration, testing, respectively. A prediction model is trained to predict the future dynamics $\mathbf{Y}\in\mathbb{R}^{d\times T}$ given the past observations $\mathbf{X}\in\mathbb{R}^{d\times\tau}$ of the particle simulation with $\tau=35,\ T=25,\ d=2$. Then all UQ methods are evaluated according to the procedures stated in Appendix IX. For UQ methods that require a further split of calibration dataset, the split ratio is set to be $0.5/0.5$ for $D_{\mathrm{cal},1}$ & $D_{\mathrm{cal},2}$ except LCP. We note that the optimization program in LCP becomes computationally intractable for this split, so we adopt $0.1/0.9$ split-ratio for LCP.
Drone trajectory
The drone trajectory dataset is generated from [sakai2018pythonroboticspythoncodecollection] with added Gaussian noise of $\sigma=0.02$. The data-split is $600/200/200$ for training/calibration/testing. The prediction model forecasts a drone’s future trajectory given its past observations ($\tau=60,\ T=10,\ d=3$). After training the prediction model, all UQ methods are evaluated in a similar manner. For methods requiring a split of the calibration dataset, a split-ratio of $0.5/0.5$ is adopted (including LCP).
IV-B Real Dataset: Covid-19 Daily Cases
We also conduct a case study on the real dataset, UK Covid-19 Daily Cases. We use the preprocessed dataset from [sun2024copula]. [^2] The original Covid-19 dataset can be download at https://coronavirus.data.gov.uk/ Each time-series data in the preprocessed Covid-19 dataset corresponds to $150$-day daily cases from mid-September $2020$ to mid-February $2021$ at a region in UK. Then, $500/160/80$ time-series data are used for training/calibration/testing. The prediction model takes $100$ days of data as input, and outputs the subsequent $50$ days, i.e., $\tau=100,T=50,d=1$. For UQ methods that require a further split of the calibration data, the split ratio was set to $0.5/0.5$.
IV-C Numerical Results
Figure 2: Performance visualization on the Particle Dataset ($\sigma=0.05$). The dashed reference line in the Coverage graph denotes the target confidences, and only methods with coverage curves at or above this line achieve the target coverages. In the Volume graph, curves closer to the bottom indicate better performances (less conservative).
For target confidence levels from 0.5 to 0.95 (10 values), we tested each UQ method with 50 runs (random splits of calibration and test set but with the same proportion) on each dataset. OSCP aims at producing minimal radius norm-balls; it is thus not direct how to compare that radius with some UQ methods whose outcome is not a norm-ball one. Thus, we consider comparing the total volume (area/length depending on the dimension) of confidence sets. A more detailed description of the comparison setup & performance evaluation of all methods is provided in Appendix IX-A.
Our method outperforms all the baselines on 4 case studies for all 10 confidence levels ($0.5$ to $0.95$). Specifically, compared to previous SOTA method LCP [cleaveland2024conformal] (the one with the smallest total volume among baselines that have empirical coverage no smaller than target ones), our method with $l_{2}$-norm, OSCP-$l_{2}$, reduces the total confidence region size (on average) of 16.03%, 14.32%, 14.01%, 16.93% on dataset Particle ($\sigma=0.01$), Particle ($\sigma=0.05$), Drone, Covid-19, respectively. When using an ellipsoidal confidence region, our method achieves further reductions in the region size (see Appendix IX-B).
Part of the experiment results (with standard error) are shown in Figure 2, and the complete results & visualizations are in Appendix IX-B. From the results in Figure 2, it can be seen that our method returns a confidence region with the smallest volume among all alternatives, while achieving the target confidence levels.
Runtime comparison
We also compare the runtime used in solving the optimization problem between our method and the previous SOTA (LCP, [cleaveland2024conformal]). For the Particle ($\sigma=0.05$) dataset, LCP sometimes reach the time limit and terminate the simulation at $10000$s, so the actual computing time is higher than the reported result. In Table I, we can see that when target confidence is set as $0.9$, our method is 8812.0, 78622.0, 14.4, 22.1 times faster than LCP on the four datasets, respectively.
TABLE I: Comparison of optimization runtime (in sec) for target confidence $1-\epsilon=0.9$
Case Study Previous SOTA [cleaveland2024conformal] OSCP-$l_{2}$
Particle ($\sigma=0.01$) 1215.002 $\pm$ 1450.119 0.137 $\pm$ 0.057
Particle ($\sigma=0.05$) >>9392.66 $\pm$ 1358.109 0.119 $\pm$ 0.034
Drone 0.151 $\pm$ 0.033 0.010 $\pm$ 0.007
Covid-19 0.549 $\pm$ 0.166 0.025 $\pm$ 0.011
V Conclusion
In this work, we propose a new parameterized score function for conformal prediction in multi-dimensional time series and an optimization program to determine an optimal parameter set. We prove validity and efficiency of our method, showing that optimizing these parameters is equivalent to determining the minimum-average-radius CP regions with a pre-specified norm-ball description. Numerical results on four different datasets (synthetic and actual data) demonstrate that our method outperforms alternative approaches, while having much lower computational requirements.
APPENDIX
VI Proof of Theorem 2
Lemma 1.
Fix an error level $\epsilon\in(0,1)$, and a predefined score function $\mathcal{A}$. Fix also any norm, and let $r_{t}^{*}$, $t=0,\ldots,T-1$, be the optimal parameters of OSCP’s score function. Any valid CP region (with same shape as OSCP’s) constructed based on $D_{\mathrm{cal},1}$ has average radius $\frac{1}{T}\sum_{t=0}^{T-1}r_{t}\geq\frac{1}{T}\sum_{t=0}^{T-1}r_{t}^{*}$.
Proof of Lemma 1: We first show that if the set (not necessarily a CP-region) $\Gamma^{\epsilon}(\hat{\mathbf{Y}}_{0:T-1}):={\otimes}^{T-1}_{t=0}\left\{y\in\mathbb{R}^{d}:||y-\hat{Y}_{t}||\leq r_{t}\right\}$ contains at least $p_{1}$ elements out of $\mathbf{Y}^{(i)}$, $i=1,\ldots,n_{1}$, then it has average radius no smaller than $\frac{1}{T}\sum_{t=0}^{T-1}r_{t}^{*}$.
Let $p_{1}=\lceil(1-\epsilon)(n_{1}+1)\rceil$. Consider the set
$\Gamma^{\epsilon}(\hat{\mathbf{Y}}_{0:T-1})^{*}:={\otimes}^{T-1}_{t=0}\left\{y\in\mathbb{R}^{d}:||y-\hat{Y}_{t}||\leq r_{t}^{*}\right\},$ (17)
where $\{r_{t}^{*}\}_{t=0}^{T-1}$ is the optimal solution to the (MILP). Since $r_{0}^{*},r_{2}^{*},...,r_{T-1}^{*}$} is feasible to (MILP), there are at least $p$ indices from $\{1,...,n_{1}\}$ such that $\tilde{e}^{(i)}_{t}\leq r_{t}^{*},\forall t=0,2,...,T-1$. Now, since $\{r_{t}^{*}\}$ is optimal to the objective of (MILP), then the average radius of $\Gamma^{\epsilon}(\hat{\mathbf{Y}}_{0:T-1})^{*}$ is the minimum among all norm-ball regions that contains at least $p_{1}$ elements out of $\mathbf{Y}^{(i)}$, $i=1,\ldots,n_{1}$.
Now, consider CP regions constructed on $D_{\mathrm{cal},1}$ via the score function $\mathcal{A}$ and a selected shape induced by $||\cdot||$:
$\Gamma^{\epsilon}(\hat{\mathbf{Y}}_{0:T-1}):=\{\mathbf{Y}\in\mathbb{R}^{T\times d}:\mathcal{A}(\mathbf{Y}-\hat{\mathbf{Y}}_{0:T-1})\leq R^{[p_{1}]}\}.$
Since $R^{[p_{1}]}$ is the $p_{1}$th smallest nonconformity score, then there are at least $p_{1}$ number of i’s satisfying
$\mathcal{A}(\tilde{\mathbf{Y}}^{(i)}_{0:T-1})\leq R^{[p_{1}]}\Rightarrow\mathbf{Y}^{(i)}_{0:T-1}\in\Gamma^{\epsilon}(\hat{\mathbf{Y}}^{(i)}_{0:T-1})$.
Then the average radius of $\Gamma^{\epsilon}(\hat{\mathbf{Y}}_{0:T-1})$ cannot be smaller than that of (17), which is $\frac{1}{T}\sum_{t=0}^{T-1}r_{t}^{*}$.  ∎
Proof of Theorem 2: Now with Lemma 1, we start proofing the Theorem 2. Given an optimal solution $\{r_{0}^{*},...,r_{T-1}^{*},b_{1}^{*},...,b_{n_{1}}^{*}\}$ of (MILP), the non-conformity score $R_{i}$ of each data is then calculated by:
$R_{i}:=\mathcal{A}(\tilde{\mathbf{Y}}^{(i)}_{0:T-1})=\max\{\tilde{e}_{0}^{(i)}-r_{0}^{*},...,\tilde{e}_{T-1}^{(i)}-r_{T-1}^{*}\}.$
To prove the result stated in the theorem, we first show that $R^{[p_{1}]}=0$. Let $i_{1},i_{2},...,i_{p_{1}}$ be the indices such that $b_{i}^{*}=1$. Due to feasibility of $r_{t}^{*}$ for any t=0,,T1t=0,...,T-1, we have $\tilde{e}_{t}^{(i)}\leq r_{t}^{*},\forall i=i_{1},...,i_{p_{1}}$. Thus, for $i=i_{1},...,i_{p_{1}}$, we have $R_{i}\leq 0$. We then have $R^{[p_{1}]}\leq\underset{i=i_{1},...,i_{p_{1}}}{\max}R_{i}\leq 0$. Now suppose for the sake of contradiction that $R^{[p_{1}]}<0$. Then $\exists i_{1}^{\prime},i_{2}^{\prime},...,i_{p_{1}}^{\prime}$ such that $R_{i}<0,\forall i=i_{1}^{\prime},i_{2}^{\prime},...,i_{p_{1}}^{\prime}$. Consequently, we have $\tilde{e}_{t}^{(i)}<r_{t}^{*},\forall t=0,...,T-1,\forall i=i_{1}^{\prime},...,i_{p_{1}}^{\prime}$.
Consider a new solution candidate
$r_{t}^{\prime}=\underset{i=i_{1}^{\prime},...,i_{p_{1}}^{\prime}}{\max}\tilde{e}_{t}^{(i)},\quad b_{i}^{\prime}=\begin{cases}1,&\text{if }i=i_{1}^{\prime},...,i_{p_{1}}^{\prime};\\ 0,&\text{otherwise}.\end{cases}$
It is easy to check that this new solution is feasible to (MILP) and $r_{t}^{\prime}<r_{t}^{*},\forall t=0,...,T-1$. This contradicts to the fact that $\sum_{t=0}^{T-1}r_{t}^{*}$ is the minimum cost solution.
Thus, we can conclude that $R^{[p_{1}]}=0$. Then for each t, the resulting CP-region is
$\{y\in\mathbb{R}^{d}$ $:||y-\hat{Y}_{t}||\leq R^{[p_{1}]}+r_{t}^{*}\}$
[…]
This completes the proof that the average radius of empirical CP-region of OSCP calculated from $D_{\mathrm{cal},1}$ is equal to $\frac{1}{T}\sum_{t=0}^{T-1}r_{t}^{*}$, which is the minimum value that a valid CP-region with same shape can attain by Lemma 1.  ∎
VII Proof of Theorem 3
Proof: Case 1: $|S_{1}|<p_{1}$
Feasibility of (MILP-fast) is easy to check, as we can always randomly pick $p_{1}$ indices $i_{1},...,i_{p_{1}}\in S$, and then the solution $\{r_{t}=\underset{i}{\max}\ \tilde{e}_{t}^{(i)}\}_{t=0}^{T-1}$, $b_{i}=\begin{cases}1,&\forall i=i_{1},...,i_{p_{1}}\\ 0,&\mathrm{otherwise}\end{cases}$ is trivially feasible to the problem. We will now prove each direction seperately.
  • ($\Rightarrow$)
    w.t.s. Any optimal parameters $\mathbf{r}^{*}=(r_{0}^{*},...,r_{T-1}^{*})$ of (MILP), $\exists\mathbf{b}^{\prime}=\{b_{i}^{\prime},\ i\in S\}$ s.t. $(\mathbf{r}^{*},\mathbf{b}^{\prime})$ is an optimal solution to (MILP-fast).
We will first show that the feasibility region of (MILP-fast) is a subset of that of (MILP).
First of all, we augment the space of decision variables $(\mathbf{r},\mathbf{b})=\{r_{0},...,r_{T-1}\}\cup\{b_{i}\}_{i\in S}$ to $(\mathbf{r},\mathbf{\bar{b}})=\{r_{0},...,r_{T-1},b_{1},...,b_{n_{1}}\}$, i.e. the feasibility region of (MILP-fast) now becomes (10), (13), (14) & (15).
Consider following constraints:
$b_{i}=1,\quad i\in S_{1};$ (18)
$b_{i}=0,\quad i\in S_{2}.$ (19)
We add these two constraints on $(\mathbf{r},\mathbf{\bar{b}})$, which has no effect on the feasible region of original decision variables $(\mathbf{r},\mathbf{b})$. That being saying, the feasibility region of (MILP) is equivalent with that of augmented (MILP-fast), i.e.,
[…]
This result is easy to check: for any solution $r_{t},b_{i}$ satisfying (18) & (19),
$b_{i}\cdot(\tilde{e}_{t}^{(i)}-r_{t})\leq 0,\ t=0,\dots,T-1,\ i=1,\dots,n_{1}$
$\Leftrightarrow b_{i}\cdot(\tilde{e}_{t}^{(i)}-r_{t})\leq 0,\ t=0,\dots,T-1,\ i\in\{1,\dots,n_{1}\}\setminus S_{2}$
$\Leftrightarrow\begin{cases}b_{i}\cdot(\tilde{e}_{t}^{(i)}-r_{t})\leq 0,\ t=0,\dots,T-1,\ i\in S,\\ \underset{i\in S_{1}}{\max}\{\tilde{e}_{t}^{(i)}\}\leq r_{t},\ t=0,\dots,T-1.\end{cases}$
Therefore, under (18) & (19), Constraint (8) $\Leftrightarrow$ Constraints (13), (14). Also, under (18) & (19),
$\sum_{i=1}^{n_{1}}b_{i}=p_{1}\Leftrightarrow\sum_{i\in S}b_{i}=p_{1}-|S_{1}|,$
hence Constraint (9) $\Leftrightarrow$ Constraint (15), and thus, we can conclude the result in (20). Consequently, we have shown that the feasible region of augmented (MILP-fast) is a subset of that of (MILP), which means the optimal objective value of (MILP-fast) is no smaller than that of (MILP).
Now we will show that the optimal objective are indeed equal by showing that any set of optimal parameters $\{r_{0}^{*},...,r_{T-1}^{*}\}$ of (MILP) is feasible in (MILP-fast).
Suppose $(\mathbf{r}^{*},\mathbf{b}^{*})=(r_{0}^{*},...,r_{T-1}^{*},b_{1}^{*},...,b_{n_{1}}^{*})$ is an optimal solution of (MILP).
First, we show that $b_{i}^{*}=0$ for $i\in S_{2}$. For $\forall i\in S_{2}$, $\tilde{e}_{t}^{(i)}>r_{t}^{(\mathrm{feas})}\text{for}\ \forall t=0,...,T-1$. Since $r_{t}^{*}$ is optimal, $\sum_{t=0}^{T-1}r_{t}^{*}\leq\sum_{t=0}^{T-1}r_{t}^{(\mathrm{feas})}$. Then there must be at least one $\hat{t}$ s.t. $r_{\hat{t}}^{*}\leq r_{\hat{t}}^{(\mathrm{feas})}$. This means that $\tilde{e}_{\hat{t}}^{(i)}>r_{\hat{t}}^{(\mathrm{feas})}\geq r_{\hat{t}}^{*}$ Thus, to satisfy the constraint (8) in (MILP), it must be that $b_{i}^{*}=0$ for $\forall i\in S_{2}$.
Next, we will show that $\exists\{b_{i}^{\prime}\}_{i\in S}$ s.t. $(\mathbf{r}^{*},\mathbf{b}^{\prime})$ is feasible to (MILP-fast). Let $i_{1},...,i_{p_{1}}$ be the indices s.t. $b_{i}^{*}=1$. Then by the above result we know that $i_{1},...,i_{p_{1}}\in\{1,...,n_{1}\}\setminus S_{2}=S\cup S_{1}$. Select $p_{1}-|S_{1}|$ indices from $\{i_{1},...,i_{p_{1}}\}\setminus S_{1}$ (this is always possible as $|S_{1}|<p_{1}$) and set $b_{i}^{\prime}=1$ for these indices, and set $b_{i}^{\prime}=0$ for the rest of the indices in $S$.
Then we have $p_{1}-|S_{1}|$ indices i’s s.t. $i\in S$ and $\tilde{e}_{t}^{(i)}\leq r_{t}^{*}$ at each t. So constraint (13) & (15) are satisfied. Now, since for $i_{1},...,i_{p_{1}}$, $b_{i}^{*}=1$, it means that we have $p_{1}$ indices i’s s.t. $\tilde{e}_{t}^{(i)}\leq r_{t}^{*}$ at each t. Recall that $\tilde{e}_{t}^{[p_{1}]}$ is the $p_{1}$th smallest element in the sorted non-descending sequence $\{\tilde{e}_{t}^{(i)}\}_{i=1}^{n_{1}}$. This means that $\tilde{e}_{t}^{[p_{1}]}\leq r_{t}^{*},\quad\forall t=0,...,T-1.$ Consequently, for $\forall i\in S_{1}$,
$\tilde{e}_{t}^{(i)}\leq\tilde{e}_{t}^{[p]}\leq r_{t}^{*},\ \forall t=0,...,T-1\ \Leftrightarrow\underset{i\in S_{1}}{\max}\{\tilde{e}_{t}^{(i)}\}\leq r_{t}^{*},$
so constraint (14) is also satisfied and we can therefore conclude that $(\mathbf{r}^{*},\mathbf{b}^{\prime})$ is feasible to (MILP-fast). As optimal value of (MILP) is always larger or equal to that of (MILP-fast), we can conclude that $(\mathbf{r}^{*},\mathbf{b}^{\prime})$ is optimal solution to (MILP-fast) and the optimal value of (MILP) and (MILP-fast) are indeed equal.
  • ($\Leftarrow$)
    w.t.s. For any optimal parameters $\mathbf{r}^{*}=(r_{0}^{*},...,r_{T-1}^{*})$ of (MILP-fast), $\exists\mathbf{b}^{\prime}=(b_{1}^{\prime},b_{2}^{\prime},...,b_{n_{1}}^{\prime})$ s.t. $(\mathbf{r}^{*},\mathbf{b}^{\prime})$ is an optimal solution to (MILP).
Suppose $(\mathbf{r}^{*},\mathbf{b}^{*})$ is an optimal solution of (MILP-fast). Consider solution $(\mathbf{r}^{*},\mathbf{b}^{\prime})=(r_{0}^{*},...,r_{T-1}^{*},b_{1}^{\prime},...,b_{n_{1}}^{\prime})$, where
$b_{i}^{\prime}=\begin{cases}b_{i}^{*},&i\in S;\\ 1,&i\in S_{1};\\ 0,&i\in S_{2}.\end{cases}$
Then we have
[…]
so constraint (9) is satisfied. For constraint (8), let’s first consider the case $i\in S\cup S_{2}$, the constraints $b_{i}^{\prime}\cdot(\tilde{e}_{t}^{(i)}-r_{t}^{*})\leq 0,\ t=0,\dots,T-1$ are trivially satisfied. For case of $i\in S_{1}$, constraint (14) in (MILP-fast) says that $b_{i}^{\prime}\cdot(\tilde{e}_{t}^{(i)}-r_{t}^{*})=\tilde{e}_{t}^{(i)}-r_{t}^{*}\leq 0,\ t=0,...,T-1$. Combining these results, constraint (8) is satisfied. As a result, the solution $(\mathbf{r}^{*},\mathbf{b}^{\prime})$ is feasible to (MILP). Since in the proof of $(\Rightarrow)$ we have already shown that the optimal value of (MILP-fast) is equal to (MILP), we can conclude that $(\mathbf{r}^{*},\mathbf{b}^{\prime})$ is optimal to (MILP).
Case 2: $|S_{1}|\geq p_{1}$
Consider the solution candidate $r_{t}^{*}=\tilde{e}_{t}^{[p_{1}]},t=0,...T-1$. We will first show it is feasible to (MILP). Pick arbitrary $p_{1}$ indices $i_{1},i_{2},...,i_{p_{1}}$ from $S_{1}$, then let $b_{i}=1,\forall i=i_{1},...,i_{p_{1}}$ and let $b_{i}=0$ otherwise. This guarantees constraints (9) & (10) satisfied. Since $\forall i\in S_{1}$, $\tilde{e}_{t}^{(i)}\leq\tilde{e}_{t}^{[p_{1}]}$ for $\forall t=0,...,T-1$, constraint (8) is also satisfied. Thus, the solution $r_{t}^{*}=\tilde{e}_{t}^{[p_{1}]},t=0,...,T-1$ is feasible to (MILP).
Now, we will show that it is also optimal to (MILP). Suppose, for contradiction, $\exists$ a feasible solution $\{r_{0}^{\prime},...,r_{T-1}^{\prime}\}$ of (MILP) such that the objective value $\sum_{t=0}^{T-1}r_{t}^{\prime}<\sum_{t=0}^{T-1}r_{t}^{*}$. Then $\exists r_{\hat{t}}^{\prime}<r_{\hat{t}}^{*}=\tilde{e}_{\hat{t}}^{[p_{1}]}$ for some $\hat{t}$. This means there are less than $p_{1}$ i’s s.t. $\tilde{e}_{\hat{t}}^{(i)}\leq r_{\hat{t}}^{\prime}$. However, constraints (8) & (9) together imply that there $\exists$ at least $p_{1}$ i’s such that $\tilde{e}_{\hat{t}}^{(i)}\leq r_{\hat{t}}^{\prime},\forall t=0,...,T-1$. Contradiction! Thus, the solution $r_{t}^{*}=\tilde{e}_{t}^{[p_{1}]},t=0,...,T-1$ is optimal to (MILP).  ∎
VIII Pseudo-code of the Fast Algorithm for Finding Optimal $\{r_{0}^{*},...,r_{T-1}^{*}\}$
Below is the pseudo-code of the faster algorithm that computes the optimal parameters $\{r_{0}^{*},...,r_{T-1}^{*}\}$ for the parameterized score function stated at the end of Section III-C.
Algorithm 1 Faster MILP for computing parameters $r_{t}$
1:Input: $\{\{\tilde{e}_{t}^{(i)}\}_{t=0}^{T-1}\}_{i=1}^{n_{1}}$, $p_{1}$
2:Output: CP parameters $r_{0}^{*},...,r_{T-1}^{*}$ // Heuristic of constructing a feasible parameter $\{r_{0}^{\mathrm{(feas)}},...,r_{T-1}^{(\mathrm{feas})}\}$
3:for $i=1,2,...,n_{1}$ do
4:  $\mathrm{ResSum}(i)\leftarrow\sum_{t=0}^{T-1}\tilde{e}_{t}^{(i)}$
5:end for
6: Sort $\mathrm{ResSum}(1),...,\mathrm{ResSum}(n_{1})$ in non-descending order
[…]
8: let $r_{t}^{(\mathrm{feas})}=\underset{i=i_{1},...,i_{p_{1}}}{\max}\tilde{e}_{t}^{(i)}$ // Finding $S_{1}$ & $S_{2}$
9:for $i=1,2,...,n_{1}$ do
10:  if $|S_{1}|=p_{1}$ then
11:   Exit and output $r_{t}=\tilde{e}_{t}^{[p_{1}]},\forall t$
12:  else if $\tilde{e}_{t}^{(i)}\leq\tilde{e}_{t}^{[p_{1}]}$ for all t=0,,T1t=0,...,T-1 then
13:   add i into $S_{1}$
14:  else if $\tilde{e}_{t}^{(i)}>r_{t}^{(\mathrm{feas})}$ for all t=0,,T1t=0,...,T-1 then
15:   add i into $S_{2}$
16:  end if
17:end for
18: Solve (MILP-fast) & output the optimal $\{r_{0}^{*},...,r_{T-1}^{*}\}$
IX Additional Details of Numerical Experiments
This section presents additional details of the numerical experiments in Section IV.
IX-A Additional Information of Experiment Setup and Evaluation Metrics
Experiment Setup
All experiments are conducted on Windows 11 (64-bit) machine with i9-13900HX CPU (24 physical cores), 32GB RAM. The computations (including prediction models training) use CPU only. For solving the optimization problems in our method and in [cleaveland2024conformal], we use the commercial solver Gurobi Optimizer (version 12.0.0 build v12.0.0rc1).
Performance Evaluation
We mainly focus on testing validity and efficiency for the UQ methods. In each case study, the dataset is first split into two halves. The first half is used for training the time-series prediction model $f_{\mathrm{pred}}$. To eliminate the variance of performance caused by model’s variance, the model is fixed once trained. Then we repeat the following procedures for 50 runs:
  1. 1.
    Randomly split the second half into calibration-set and test-set with fixed proportion
  2. 2.
    For error tolerance ($\epsilon$) from 0.05 to 0.5 (10 different values in total), use calibration-set to train the UQ methods.
  3. 3.
    Test the UQ methods on test-set and calculate the validity & efficiency metrics (see below) on test-set for $\epsilon=0.05$ to $0.5$.
Lastly, all the results are averaged over the 50 runs.
Validity Metric
At each run, the validity is evaluated by the empirical coverage on the test-set, which is calculated as follows:
$\text{Coverage}(\Gamma^{\epsilon}):=\frac{1}{|D_{test}|}\sum_{(\mathbf{X}^{(i)},\mathbf{Y}^{(i)})\in D_{test}}\mathbb{I}\left(\mathbf{Y}^{(i)}\in\Gamma^{\epsilon}(\mathbf{X}^{(i)})\right).$ (21)
Efficiency Metric
Since the width of the confidence set is difficult to calculate for some UQ methods, we instead consider calculating the total volume (Lebesgue measure) of the confidence sets as efficiency metric $\mathcal{L}_{\mathrm{eff}}$ of UQ methods. For $d=1$ & $d=2$ cases, the "volume" corresponds to length and area, respectively. For an error tolerance $\epsilon$, the total volume of a UQ method is computed as follows:
$\text{Volume}(\Gamma^{\epsilon}):=\frac{1}{|D_{test}|}\sum_{\mathbf{X}^{(i)}\in D_{test}}Volume\left(\Gamma^{\epsilon}(\mathbf{X}^{(i)})\right).$ (22)
Training Details of time-series Forecasting Model $f_{\mathrm{pred}}$
Since we use the similar experiment settings (e.g., datasets, evaluation metrics, etc.) as in [sun2024copula], we also use the experiment code of [sun2024copula] to train the underlying time-series prediction model for all 4 case studies. [^3] Experiment code of [sun2024copula] can be download at https://github.com/Rose-STL-Lab/CopulaCPTS Here is a brief description of the forecasting model $f_{\mathrm{pred}}$ for 4 case studies.
The forecasting model for Particle Datasets ($\sigma=0.01$ &\& $\sigma=0.05$) is an RNN network with embedding size == 24, where the hidden state is then passed through a linear network to concurrently predict the time-steps. Then the model is trained for 150 epochs and batch size == 150.
The forecasting model for the Drone Dataset is an RNN with embedding size == 128 that is trained with 500 epochs and batch size == 150.
For the Covid-19 Dataset, the forecasting model is an RNN with embedding size==128, and is trained with epochs == 200 & batch size == 50.
IX-B Detailed Experiment Results and Visualizations
The complete results of our method with $l_{2}$-norm (OSCP-$l_{2}$) & with ellipsoidal norm (OSCP-Ell) and of all 5 baselines (LCP [cleaveland2024conformal], CopulaCPTS [sun2024copula], CRD [Tumu2024multi-modal], CF-RNN [stankeviciute2021conformal], MC-dropout [gal2016dropout]) on all 4 case studies are shown in Table II, Figure 3 & Figure 4.
In Table II, the data in bold have better efficiency (smaller total volume) than previous SOTA (LCP, [cleaveland2024conformal]) and have coverages higher or equal to the target confidences, while the data in gray mean the empirical coverages are lower than the target confidences. It is important to note that if the empirical coverage of a method is slightly lower than the target confidence level (e.g. difference of coverages $\leq 2\%$), it does NOT necessarily mean that this method is invalid. The small gap in coverage can also be caused by numerical inaccuracy in computation (e.g., matrix inverse, eigen-value computations), or simply the randomness of empirical coverage (which converges to $\lceil(1-\epsilon)\cdot(N_{cal,2}+1)\rceil/(N_{cal,2}+1)$ in probability as $|D_{test}|\rightarrow\infty$, by the weak law of large numbers). However, to ensure a fair comparison of efficiency, here we only compare the results whose empirical coverages are no smaller than the target ones.
Figure 3: Case studies: Particle Datasets. The dashed reference line denotes the target confidences, and only methods with coverage curves at or above this line achieve the target coverages. The shaded region of each curve is the $\pm$ 1 standard error region. In Area graphs, lower curves indicate better performance.
Figure 4: Case studies: Drone Dataset & Covid-19 Dataset. The dashed reference line denotes the target confidences, and only methods with coverage curves at or above this line achieve the target coverages. The shaded region of each curve is the $\pm$ 1 standard error region. In Length/Volume graphs, lower curves indicate better performance.
TABLE II: Detailed performance comparison of all baselines with confidence levels from 85% to 95%
(a) Particle trajectory simulation ($\sigma=0.01$): $d=2$, $T=25$
Target Coverage: 85% Target Coverage: 90% Target Coverage: 95%
Method Coverage Area Coverage Area Coverage Area
OSCP-$l_{2}$ (Ours) 85.8 $\pm$ 2.7 0.59 $\pm$ 0.03 90.5 $\pm$ 2.1 0.69 $\pm$ 0.04 95.4 $\pm$ 1.6 0.88 $\pm$ 0.07
OSCP-EllEll (Ours) 85.3 $\pm$ 2.8 0.59 $\pm$ 0.04 90.2 $\pm$ 2.4 0.68 $\pm$ 0.05 95.3 $\pm$ 1.7 0.87 $\pm$ 0.07
LCP [cleaveland2024conformal] 85.4 $\pm$ 2.2 0.69 $\pm$ 0.07 90.1 $\pm$ 2.3 0.81 $\pm$ 0.08 95.5 $\pm$ 1.4 1.05 $\pm$ 0.12
CopulaCPTS [sun2024copula] 85.3 $\pm$ 3.3 0.74 $\pm$ 0.11 89.8 $\pm$ 3.0 0.91 $\pm$ 0.3 94.4 $\pm$ 2.1 1.7 $\pm$ 1.16
CRD-Ell [Tumu2024multi-modal] 85.3 $\pm$ 3.0 0.93 $\pm$ 0.23 90.2 $\pm$ 2.4 1.02 $\pm$ 0.22 95.6 $\pm$ 1.4 1.24 $\pm$ 0.19
CRD-Rect [Tumu2024multi-modal] 92.7 $\pm$ 4.2 1.52 $\pm$ 0.41 94.5 $\pm$ 3.3 1.55 $\pm$ 0.41 96.1 $\pm$ 2.2 1.59 $\pm$ 0.39
CF-RNN [stankeviciute2021conformal] 97.0 $\pm$ 1.7 1.19 $\pm$ 0.22 98.3 $\pm$ 1.5 2.04 $\pm$ 1.03 99.4 $\pm$ 0.7 4.12 $\pm$ 1.46
[…]
(b) Particle trajectory simulation ($\sigma=0.05$): $d=2$, $T=25$
Target Coverage: 85% Target Coverage: 90% Target Coverage: 95%
Method Coverage Area Coverage Area Coverage Area
OSCP-$l_{2}$ (Ours) 85.4 $\pm$ 2.4 3.85 $\pm$ 0.18 90.6 $\pm$ 2.4 4.43 $\pm$ 0.27 95.6 $\pm$ 1.7 5.6 $\pm$ 0.4
OSCP-EllEll (Ours) 86.1 $\pm$ 2.5 3.81 $\pm$ 0.16 90.5 $\pm$ 2.2 4.27 $\pm$ 0.17 95.2 $\pm$ 1.7 5.16 $\pm$ 0.31
LCP [cleaveland2024conformal] 85.4 $\pm$ 2.4 4.43 $\pm$ 0.29 90.5 $\pm$ 1.8 5.07 $\pm$ 0.3 95.5 $\pm$ 1.4 6.41 $\pm$ 0.67
CopulaCPTS [sun2024copula] 85.6 $\pm$ 2.8 4.44 $\pm$ 0.46 90.0 $\pm$ 2.6 5.43 $\pm$ 0.86 93.4 $\pm$ 2.2 6.46 $\pm$ 1.28
CRD-Ell [Tumu2024multi-modal] 85.6 $\pm$ 2.8 4.88 $\pm$ 0.44 90.5 $\pm$ 2.0 5.5 $\pm$ 0.58 95.6 $\pm$ 1.1 7.11 $\pm$ 0.71
CRD-Rect [Tumu2024multi-modal] 92.0 $\pm$ 3.0 7.09 $\pm$ 0.82 93.2 $\pm$ 2.4 7.3 $\pm$ 1.05 95.4 $\pm$ 1.8 7.9 $\pm$ 1.06
CF-RNN [stankeviciute2021conformal] 92.6 $\pm$ 1.5 5.2 $\pm$ 0.29 95.0 $\pm$ 1.5 6.08 $\pm$ 0.67 97.6 $\pm$ 1.1 8.41 $\pm$ 1.29
MC-dropout [gal2016dropout] 33.8 $\pm$ 4.8 1.81 $\pm$ 0.07 42.7 $\pm$ 5.1 2.36 $\pm$ 0.09 55.5 $\pm$ 5.2 3.35 $\pm$ 0.13
(c) Drone trajectory simulation ($\sigma=0.02$): $d=3$, $T=10$
Target Coverage: 85% Target Coverage: 90% Target Coverage: 95%
Method Coverage Volume Coverage Volume Coverage Volume
OSCP-$l_{2}$ (Ours) 85.2 $\pm$ 4.6 7.58 $\pm$ 1.09 90.2 $\pm$ 3.7 8.59 $\pm$ 1.43 95.5 $\pm$ 2.8 11.22 $\pm$ 2.84
OSCP-EllEll (Ours) 83.6 $\pm$ 3.9 5.3 $\pm$ 1.25 88.6 $\pm$ 4.2 7.11 $\pm$ 2.4 94.8 $\pm$ 2.3 11.01 $\pm$ 4.47
LCP [cleaveland2024conformal] 84.4 $\pm$ 4.5 8.51 $\pm$ 2.09 89.9 $\pm$ 3.2 9.2 $\pm$ 2.29 95.2 $\pm$ 2.9 11.91 $\pm$ 4.83
CopulaCPTS [sun2024copula] 87.8 $\pm$ 5.0 9.94 $\pm$ 1.92 91.5 $\pm$ 4.2 10.79 $\pm$ 2.07 95.6 $\pm$ 2.6 12.74 $\pm$ 1.95
CRD-Ell [Tumu2024multi-modal] 84.7 $\pm$ 3.7 163.64 $\pm$ 161.07 90.0 $\pm$ 3.1 207.23 $\pm$ 221.37 95.1 $\pm$ 2.1 305.0 $\pm$ 537.19
[…]
MC-dropout [gal2016dropout] 75.0 $\pm$ 11.6 4.5 $\pm$ 2.1 82.6 $\pm$ 10.5 6.76 $\pm$ 3.22 90.3 $\pm$ 8.0 11.35 $\pm$ 5.22
(d) Covid-19 daily cases: $d=1$, $T=50$
Target Coverage: 85% Target Coverage: 90% Target Coverage: 95%
Method Coverage Length Coverage Length Coverage Length
OSCP-$l_{2}$ (Ours) 86.2 $\pm$ 5.6 106.01 $\pm$ 11.72 91.4 $\pm$ 4.8 133.4 $\pm$ 19.5 96.4 $\pm$ 3.6 186.23 $\pm$ 29.83
[…]
CopulaCPTS [sun2024copula] 73.2 $\pm$ 8.0 93.45 $\pm$ 17.03 77.1 $\pm$ 6.6 104.86 $\pm$ 14.86 81.4 $\pm$ 6.3 116.53 $\pm$ 14.42
CRD-Ell [Tumu2024multi-modal] 85.9 $\pm$ 5.6 1206.44 $\pm$ 238.9 91.2 $\pm$ 4.9 1486.12 $\pm$ 365.38 96.7 $\pm$ 3.3 2074.58 $\pm$ 549.95
CRD-Rect [Tumu2024multi-modal] 85.9 $\pm$ 5.7 798.16 $\pm$ 167.12 91.3 $\pm$ 4.8 999.14 $\pm$ 277.5 77.3 $\pm$ 39.2 nan $\pm$ nan
CF-RNN [stankeviciute2021conformal] 92.8 $\pm$ 3.8 165.4 $\pm$ 10.66 92.8 $\pm$ 3.8 165.4 $\pm$ 10.66 92.8 $\pm$ 3.8 165.4 $\pm$ 10.66
MC-dropout [gal2016dropout] 12.8 $\pm$ 4.7 7.48 $\pm$ 0.61 15.6 $\pm$ 4.9 8.55 $\pm$ 0.7 19.3 $\pm$ 5.0 10.19 $\pm$ 0.83
IX-C Further Discussions on the Experiment Results
For deep UQ methods, MC-dropout [gal2016dropout] does not have validity guarantees and is thus far below the reference target levels for Particle ($\sigma=0.05$), Drone, and Covid-19 datasets; CF-RNN [stankeviciute2021conformal] is one of the classic works of CP in time series which holds the validity guarantee, but it offers overly large confidence regions. For the baselines from newer works, CRD [Tumu2024multi-modal] is unstable and produces excessive large confidence sets that cannot be shown in Figure 4. We hypothesize that this is due to the error accumulation during the fitting and estimation procedures of the CRD method. The performance of CopulaCPTS [sun2024copula] is also unstable. It has large variances for some cases (see the area graph with confidence = 85%, 90% & 95% in Figure 3), and also has unstable coverages (see results on Covid-19 dataset). Among all baselines, LCP [cleaveland2024conformal] has the best performance as it achieves target coverages while having smaller volumes than other baselines. However, the computation cost for LCP can be very large (see Table I in Section IV-C), and it still has larger volumes than our method, OSCP.
We can also observe that the size reduction is most significant on the Covid-19 dataset. This dataset has $d=1$, which means the total volume in this case is equivalent as the total width. As our method OSCP is originally built to produce minimal-total-width regions rather than minimal-total-volume ones, it explains why the volume reduction is highest on the Covid-19 dataset.
A general conclusion from the results in Table II and Figure 3 & 4 is that our method with $l_{2}$-norm, OSCP-$l_{2}$, outperforms all the baselines on 4 case studies for all 10 confidence levels ($0.5\%$ to $0.95\%$). Specifically, OSCP-$l_{2}$ has coverages higher than all the target confidence levels, and it achieves volume reductions (average value over 10 confidence levels) of 16.03%, 14.32%, 14.01%, 16.93% compared to previous SOTA (LCP, [cleaveland2024conformal]) on Particle ($\mathbf{\sigma=0.01}$), Particle ($\mathbf{\sigma=0.05}$), Drone, Covid-19 datasets, respectively.
Besides, our method with ellipsoidal norm, OSCP-$Ell$, achieves volume reduction of 13.65%, 14.81%, 33.5% compared to previous SOTA (LCP, [cleaveland2024conformal]) on Particle ($\mathbf{\sigma=0.01}$), Particle ($\mathbf{\sigma=0.05}$), Drone, respectively (since there is no ellipsoid for 1d case, we did not run OSCP-$Ell$ on the Covid-19 dataset). Ellipsoidal norm fitted an ellipsoid shape for each time step using $D_{\mathrm{cal,1}}$, and ellipsoid norm performs better when the residuals exhibit greater non-Gaussianity and/or when the dimension $d$ is larger. This explains why we can observe a more significant volume reduction on Drone dataset ($d=3$) using OSCP-$Ell$ (33.5%) compared to OSCP-$l_{2}$ (14.1%), while we cannot see significant differences on the Particle datasets (13.65% vs. 16.03% and 14.81% vs. 14.32%, respectively). In a word, when $|D_{\mathrm{cal}}|$ is small and we have no prior knowledge of the behavior of the residual at each time step, we can just assume the residual at each time step is Gaussian-error and use OSCP-$l_{2}$. Otherwise, if we know that the residual is highly non-Gaussian and the $|D_{\mathrm{cal}}|$ is not too small, using OSCP-$Ell$ is more preferable.
IX-D Construction of Ellipsoid-Norm
Given initial observations $\mathbf{X}$ of a time series, suppose $\mathbf{Y}=(Y_{1},...,Y_{T})\in\mathbb{R}^{d\times T}$ are the true future values, and $\hat{\mathbf{Y}}=(\hat{Y}_{1},...,\hat{Y}_{T})\in\mathbb{R}^{d\times T}$ are the predicted values from $f_{\mathrm{{pred}}}$. For each time step $t$, the Ellipsoid-norm of residual $Y_{t}-\hat{Y}_{t}$ is defined by
[…]
where $\hat{\Sigma}^{-1}$ is the inverse of empirical covariance matrix estimated from $Y_{t}^{(i)}-\hat{Y}_{t}^{(i)}$ for $i=1,...,n_{1}$ (i.e., data from $D_{\mathrm{cal,1}}$).
We wish to highlight that the ellipsoid norm is fitted for each time step individually, so we have $T$ different ellipsoidal norms in total, which ensures a locally-adaptive shape fitting. To guarantee robust construction of ellipsoid-norms against numerical issues, one can use pseudo-inverses of covariance matrices via the singular-value-decomposition technique instead of the standard inverses (see [Xu2024Conformal]).